The top 30% of high school graduates receive the highest TOPS award possible while the bottom 45% receive the minimum TOPS award. Assuming the GPA scores of graduating high school students are normally distributed with a mean of 2.6 and a standard deviation of .36, which of the following statements is most accurate?
Select one:
a. To be in the top 30% of TOPS recipients, a student must have achieved a high school GPA of at least 2.79.
b. To be in the bottom 45% of TOPS recipients, a student must have achieved a high school GPA no greater than 2.50.
c. To be in the top 30% of TOPS recipients, a student must have achieved a high school GPA not greater than 2.41.
d. To be in the bottom 45% of TOPS recipients, a student must have achieved a high school GPA no greater than 2.45.

Answer by ewatrrr(24785)

a. .5244*.36 + 2.6 = .1888 + 2.6 = 2.7888 = 2.79 Yes
b. -.1257*.36 + 2.6 = -.045 + 2.6 = 2.55 (NO on the 2.5)
b. NO
c. NO